Ticket Number: HEA-NOC/19990331-02 Ticket Status: CLOSE
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: HEAnet <==> Janet
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 19990331 16:50 UTC+1
Ticket Closed: 19990401 15:35:26 UTC Problem End: 19990331 18:30 UTC+1
CRC errors on the line from HEAnet to Queen's University,
Our BGP session with JANET waws hit by the high number
of errors, so traffic for JANET and TEN-155 backed up via
the US lines with impeded performance.
Reported to Janet operations desk - awaiting further news.
Update 19990331 17:55 davew
Janet (JOD3468) checked there end and believe it to be
a line fault (so do I). Reported to Telecom Eireann, ref. 1612.
Update 19990401 14:30 davew
Telecom failed to find any fault on the line. Problem had
disappeared as of 18:30 Wednesday. Allowed T.E. and Janet
to close the ticket.
Time to Fix:
1h 40m Additional Notes: